# HydroGEM: A Self-Supervised Zero Shot Hybrid TCN-Transformer Foundation Model for Continental Scale Streamflow Quality Control

### v2

@@ 3.1 Two-Stage Training Framework
The first stage trains a deep autoencoder backbone through self-supervised learning on clean hydrological sequences, allowing the model to discover core relationships between discharge, stage, basin characteristics, and temporal dynamics without requiring explicit anomaly labels. This pretraining phase processes the 6.03 million training sequences described in Section 3.2.3 to develop general-purpose hydrological representations that capture the full spectrum of normal flow behaviors across continental scales. The second stage augments this pretrained backbone with a specialized detection head and fine-tunes the system using the synthetically generated anomalies described in Section 3.3.2—approximately 11 simplified corruption patterns applied in normalized space with controlled coverage (mean 15.2% $\pm$ 3.1%).
This decoupled approach provides several advantages over end-to-end anomaly detection architectures. The self-supervised pretraining exploits massive volumes of unlabeled data that would be prohibitively expensive to manually annotate, learning subtle hydrological patterns that might not be captured from limited labeled examples. The pretrained representations show superior transfer learning capabilities [60], allowing detection of new anomaly types not explicitly included in the fine-tuning dataset. Furthermore, the modular design permits independent optimization of reconstruction quality and detection performance, facilitating ablation studies and architectural improvements without complete retraining.
@@ 3.2.1 Backbone Network Architecture
The encoder maps 12-dimensional inputs to 128-dimensional hidden representations through four stacked TCN blocks with exponentially increasing dilation rates ($r\in\{1,2,4,8\}$). Each block uses residual connections, batch normalization, and dropout ($p=0.2$). With kernel size $k=3$, this yields receptive fields of 5, 13, 29, and 61 hours for blocks 1-4, capturing hydrological processes from sub-daily fluctuations to multi-day storm events. The TCN encoder contributes 2.1 million parameters.
@@
The TCN output is projected from 128 to 256 dimensions and processed through 4 transformer layers with 8-head attention. We use Cosine Retention Attention with learnable temporal decay [36] to prevent attention collapse in long sequences. Sliding window attention (256 positions) reduces complexity from $O(T^{2})$ to $O(T\cdot W)$. The transformer contributes 9.8 million parameters.
@@
The decoder mirrors the encoder, with four TCN blocks reconstructing the original sequence. A gated skip connection adaptively combines encoder and decoder pathways, creating a curriculum learning effect: early training relies on skip connections for gradient flow while later training strengthens the transformer pathway. The gate parameter $\alpha$ is initialized to yield near-zero contribution from the skip path, gradually increasing during training.
The complete backbone comprises 14.2 million parameters: TCN encoder (2.1M), transformer (9.8M), TCN decoder (2.1M), and projection layers (0.2M).
@@ 3.2.2 Masking Strategy and Objectives
The masking strategy in Stage 1 serves a fundamentally different purpose than the anomaly injection in Stage 2 (Section 2.3). While Stage 2 trains the model to detect corrupted values, Stage 1 establishes a robust representation of normal hydrological behavior through reconstruction of missing data. This pretraining creates a strong inductive bias—the model learns the statistical regularities, physical constraints, and temporal dynamics that characterize clean hydrological data. When subsequently exposed to anomalies during Stage 2, the model can detect them precisely because they violate these learned patterns of normality.
@@
  • Point masking (40% probability): Randomly masks 15% of timesteps, forcing reconstruction from immediate temporal context. This teaches the model short-term hydrological persistence and smoothness constraints.
  • Block masking (30% probability): Masks 1-3 contiguous blocks of 12-72 hours, requiring the model to learn recession curves, diurnal patterns, and multi-day correlations essential for bridging extended gaps.
  • Periodic masking (20% probability): Masks 4-hour windows at 168-hour intervals, teaching the model to recognize weekly operational cycles and anthropogenic patterns in streamflow.
  • Feature masking (10% probability): Masks either discharge (70%) or stage (30%) for 24-168 hours, forcing the model to learn rating curve relationships [25] and cross-variable dependencies.
Each training sequence undergoes masking with probability 0.80. The resulting reconstruction task requires the model to internalize three critical aspects of hydrological data: (1) temporal dynamics including recession rates and hydrograph shapes, (2) physical relationships between discharge and stage governed by hydraulic geometry, and (3) scale-dependent patterns that vary with basin size and flow regime. These learned representations provide the foundation for detecting anomalies that violate expected patterns during operational deployment.
@@ 3.2.3 Pretraining Loss Functions
  • Diversity regularization: Prevents attention collapse through entropy maximization
The backbone trains for 20 epochs using AdamW optimization [33] with OneCycleLR scheduling (peak learning rate $5\times 10^{-4}$), gradient clipping (norm 1.0), and early stopping (patience 7 epochs). Training requires approximately 48 GPU-hours on NVIDIA A100 hardware.
3.3 Stage 2: Anomaly Detection Fine-Tuning
The second stage augments the pretrained backbone with a specialized detection head and fine-tunes the complete system using the synthetic anomaly injection framework described in Section 2.3. As detailed in that section, the training injector implements approximately 11 simplified anomaly patterns (spikes, drift, flatlines, dropouts, saturation, clock shifts, quantization, unit jumps, warping, splicing, and subtle drift) applied in normalized log-space with controlled coverage targeting 15.2% $\pm$ 3.1% temporal coverage. This deliberate simplification—using simple corruption mechanisms during training while testing on complex physical-space anomalies (Section 2.4)—forces the model to learn fundamental hydrometric consistency principles rather than memorizing specific anomaly signatures.
@@ 3.3.1 Detection Head Architecture
The detection head operates entirely on observable quantities without requiring ground-truth clean data during inference [3, 16]. It computes an 11-dimensional feature vector from the relationship between potentially corrupted observations and backbone reconstructions:
  • Reconstruction residuals: Absolute differences between observed and reconstructed discharge/stage
@@
  • Rating curve deviation: Local power-law fit quality indicating hydraulic inconsistency
@@
Features undergo robust standardization using median absolute deviation and are processed through a two-layer MLP (128 and 64 units) with GELU activation and dropout ($p=0.2$). Monte Carlo dropout (10 forward passes) provides uncertainty quantification for operational decision-making.
3.3.2 Fine-Tuning Objectives
The fine-tuning loss balances four objectives (complete formulations in Appendix C):
@@ 3.3.2 Fine-Tuning Objectives
  • Physics constraints: Enforces discharge-stage coupling and rating curve compliance
Fine-tuning proceeds for 12 epochs with AdamW ($lr=5\times 10^{-5}$), requiring approximately 8 GPU-hours.
@@ 3.4 Inference Pipeline
At deployment, the trained model processes incoming observations through the frozen backbone to generate reconstructions, then computes detection features from observation-reconstruction residuals (Figure 2, bottom). The detection head outputs anomaly probabilities with uncertainty estimates obtained via MC Dropout ($N=20$ forward passes). High-uncertainty predictions are flagged for mandatory expert review. For detected anomalies, the backbone reconstruction serves as a suggested correction—not an automated replacement—requiring hydrologist approval before integration into official records. This human-in-the-loop design ensures data quality while enabling the model to screen thousands of stations continuously.
@@ 4.1 Computational Infrastructure
All experiments were run on the Lonestar6 supercomputer at the Texas Advanced Computing Center, accessed through the NSF National AI Research Resource (NAIRR) Pilot program. The backbone pretraining phase ran for 20 epochs and required approximately 46 hours on NVIDIA A100 GPUs. The fine tuning phase completed in 12 epochs. For both stages we selected the checkpoint with the lowest validation loss and used that model state for all downstream evaluation.
@@ 4.2 Evaluation Metrics
We evaluate HydroGEM along 2 complementary dimensions: anomaly detection performance and reconstruction quality.
For anomaly detection we report precision, recall, F1 score, and the area under the receiver operating characteristic curve (AUROC). Precision measures the fraction of flagged points that are true anomalies, recall measures the fraction of true anomalies that are correctly flagged, and F1 is the harmonic mean of precision and recall. F1 is our primary reporting metric, with AUROC providing a threshold independent view of ranking quality.
@@
$\text{Error Reduction}=\frac{\lvert X^{(\text{raw})}-X^{(\text{clean})}\rvert-\lvert\hat{X}-X^{(\text{clean})}\rvert}{\lvert X^{(\text{raw})}-X^{(\text{clean})}\rvert}\times 100\%,$ (1)
and report segment wise averages together with root mean squared error (RMSE) on anomalous segments. We additionally track RMSE on non anomalous timesteps to confirm that the model does not degrade clean observations. On the synthetic experiments HydroGEM typically reduces error by around 70% on corrupted segments, while keeping errors on clean segments close to the baseline.
For the Canadian case study, reconstructions are evaluated more qualitatively. Weak labels are derived from operational corrections, and correction strategies, especially for ice affected periods where discharge is often pushed very low while stage remains relatively high, differ from the USGS corrections used to train HydroGEM. Official guidance notes that stage affected by backwater from ice may accurately represent local conditions but may not be suitable for discharge computation, and operational products explicitly warn that ice can raise stage for a given flow [41, 44, 18]. In our sample many Environment and Climate Change Canada winter corrections reduce discharge much more aggressively than the USGS style corrections seen in training, particularly during ice control periods where stage remains elevated. As a result, in the Canadian evaluation we treat reconstruction error as a secondary diagnostic rather than a primary performance metric and place emphasis on detection metrics.
@@ 4.3 Baseline Methods
To assess zero shot anomaly detection performance we compare HydroGEM against 12 baseline methods that do not require labeled training data or site specific calibration. All baselines operate in a zero shot mode at new stations, matching HydroGEM’s deployment setting. We consider 3 categories.
@@
These methods represent simple distributional checks that are widely used in practice. They include a Z Score detector that flags observations with absolute z score greater than 3 for either discharge or stage, an Interquartile Range (IQR) rule that flags values outside the interval $Q_{1}-1.5\times\text{IQR},Q_{3}+1.5\times\text{IQR}]$, and a Moving Average Residual detector that uses a centered 7 day window and flags points whose residuals exceed 3 standard deviations of the rolling window. Together these capture point wise outliers and deviations from local temporal context.
@@
We include generic unsupervised anomaly detectors that require no labels and can be applied at scale. Isolation Forest and Local Outlier Factor (LOF) operate on standardized discharge and stage and use standard contamination settings, representing tree based and density based approaches to unsupervised outlier detection. A Seasonal Trend decomposition using LOESS (STL) baseline models a 7 day seasonal component and flags residuals that exceed 3 standard deviations. These methods provide a representative sample of modern off the shelf anomaly detectors that do not use hydrological structure.
@@
The third group encodes relationships that hydrologists routinely use during manual quality control. A Rating Curve Residual baseline fits a power law relationship $Q=a(H-H_{0})^{b}$ by log-linear regression and flags points with large log-space residuals. This serves as a key reference because rating residuals are central to operational practice for detecting rating shifts, drift, and ice effects. Additional checks include rate-of-change limits on discharge and stage, persistence detection for stuck sensors, Q-H consistency checks that flag periods where discharge and stage trends diverge or correlations break down, and a seasonal envelope that flags values outside month-specific percentile bands. Together these methods approximate the mix of physical rules and heuristics that agencies use in production systems.
@@
We intentionally exclude methods that rely on extensive labeled anomalies or site specific tuning. Supervised deep learning classifiers and autoencoders require labels for each station and would not represent a true zero shot setting. Per site reconstruction models that are fit directly on test windows are also excluded, since they would not be deployed in this way operationally. Proprietary automated flagging systems from agencies such as USGS and Environment and Climate Change Canada are not reproducible without internal configuration parameters. Pretraining multiple alternative deep architectures on the full USGS corpus would be methodologically interesting but is beyond the present computational scope and is left for future work.

### v3

@@ 3.1 Two-Stage Training Framework
The first stage trains a deep autoencoder backbone through self-supervised learning on clean hydrological sequences, allowing the model to discover core relationships between discharge, stage, basin characteristics, and temporal dynamics without requiring explicit anomaly labels. This pretraining phase processes the 6.03 million training sequences described in Section 2.1.4 to develop general-purpose hydrological representations that capture the full spectrum of normal flow behaviors across continental scales. The second stage freezes the pretrained backbone and trains a lightweight detection head using the synthetically generated anomalies described in Section 2.3, which implements approximately 11 simplified corruption patterns applied in normalized space with controlled coverage (mean 15.2% $\pm$ 3.1%).
This decoupled approach provides several advantages over end-to-end anomaly detection architectures. The self-supervised pretraining exploits large volumes of unlabeled data that would be impractical to manually annotate at this scale, learning subtle hydrological patterns that might not be captured from limited labeled examples. The pretrained representations can support transfer to anomaly patterns not explicitly represented in the training injector [88]. Furthermore, the modular design permits independent optimization of reconstruction quality and detection performance, facilitating ablation studies and architectural improvements without complete retraining.
@@ 3.2.1 Backbone Network Architecture
The encoder maps 12-dimensional inputs to 128-dimensional hidden representations through four stacked TCN blocks with exponentially increasing dilation rates ($r\in\{1,2,4,8\}$). Each block uses residual connections, batch normalization, and dropout ($p=0.2$). With kernel size $k=3$, this yields receptive fields spanning roughly 5 to 61 hours across blocks, capturing hydrological processes from sub-daily fluctuations to multi-day storm events. The TCN encoder contributes 2.1 million parameters.
@@
The TCN output is projected from 128 to 256 dimensions and processed through 4 transformer layers with 8-head attention. We use Cosine Retention Attention with learnable temporal decay [48, 73] to improve stability on long sequences. Sliding window attention with window size $W=256$ reduces complexity from $O(T^{2})$ to $O(T\cdot W)$. The transformer contributes 9.8 million parameters.
@@
The decoder mirrors the encoder, with four TCN blocks reconstructing the original sequence. A gated skip connection adaptively combines encoder and decoder pathways, which can improve gradient flow during early training: the gate parameter $\alpha$ is learned during training, initialized to favor the main pathway. The complete backbone comprises 14.2 million parameters: TCN encoder (2.1M), transformer (9.8M), TCN decoder (2.1M), and projection layers (0.2M).
@@ 3.2.2 Masking Strategy and Objectives
The masking strategy in Stage 1 serves a fundamentally different purpose than the anomaly injection in Stage 2 (Section 2.3). While Stage 2 trains the model to detect corrupted values, Stage 1 establishes a robust representation of normal hydrological behavior through reconstruction of missing data. This pretraining creates a strong inductive bias: the model learns the statistical regularities, physical constraints, and temporal dynamics that characterize clean hydrological data. When subsequently exposed to anomalies during Stage 2, the model can detect them precisely because they violate these learned patterns of normality.
@@
  • Point masking (40% probability): Randomly masks 15% of timesteps, forcing reconstruction from immediate temporal context. This pattern addresses telemetry dropouts and transmission noise common in real-time data streams.
  • Block masking (30% probability): Masks 1-3 contiguous blocks of 12-72 hours, requiring the model to learn recession curves, diurnal patterns, and multi-day correlations. This pattern reflects sensor outages and extended maintenance windows.
  • Periodic masking (20% probability): Masks 4-hour windows at 168-hour intervals, teaching the model to recognize weekly operational cycles. This pattern can reflect scheduled operational patterns and anthropogenic influences where present.
  • Feature masking (10% probability): Masks either discharge (70%) or stage (30%) for 24-168 hours, forcing the model to learn rating-curve relationships [31] and cross-variable dependencies essential for detecting coupled sensor failures.
Each training sequence undergoes masking with probability 0.80. The model learns to reconstruct discharge and stage (the primary targets), while other features provide contextual information. The resulting reconstruction task requires the model to internalize three critical aspects of hydrological data: (1) temporal dynamics including recession rates and hydrograph shapes, (2) physical relationships between discharge and stage governed by hydraulic geometry, and (3) scale-dependent patterns that vary with basin size and flow regime. These learned representations provide the foundation for detecting anomalies that violate expected patterns during operational deployment.
@@ 3.2.3 Pretraining Loss Functions
  • Attention regularization: Encourages diverse attention patterns across heads
The backbone trains for 20 epochs using AdamW optimization [Loshchilov2017] with OneCycleLR scheduling (peak learning rate $5\times 10^{-4}$), gradient clipping (norm 1.0), and early stopping (patience 7 epochs). Training requires approximately 48 GPU-hours on NVIDIA A100 hardware.
3.3 Stage 2: Anomaly Detection Head Training
The second stage freezes the pretrained backbone and trains a lightweight detection head using the synthetic anomaly injection framework described in Section 2.3. As detailed in that section, the training injector implements approximately 11 simplified anomaly patterns (spikes, drift, flatlines, dropouts, saturation, clock shifts, quantization, unit jumps, warping, splicing, and subtle drift) applied in normalized log-space with controlled coverage targeting 15.2% $\pm$ 3.1% temporal coverage. This deliberate simplification forces the model to learn fundamental hydrometric consistency principles rather than memorizing specific anomaly signatures: simple corruption mechanisms are used during training while testing employs complex physical-space anomalies (Section 2.3.1).
@@ 3.3.1 Detection Head Architecture
The detection head operates entirely on observable quantities without requiring ground-truth clean data during inference [5, 18]. It computes an 11-dimensional feature vector from the relationship between potentially corrupted observations and backbone reconstructions. These features are chosen to reflect operationally observable inconsistencies between the observations and the model’s reconstruction, without requiring access to a ground-truth clean series:
  • Reconstruction residuals: Absolute differences between observed and reconstructed discharge and stage
@@
  • Rating-curve deviation: Local power-law fit quality indicating hydraulic inconsistency
@@
Features undergo robust standardization using median absolute deviation and are processed through a two-layer MLP (128 and 64 units) with GELU activation and dropout ($p=0.2$), yielding approximately 10K trainable parameters. Dropout is enabled during training; at inference we use Monte Carlo dropout with N=20 stochastic forward passes to estimate predictive uncertainty.
3.3.2 Detection Head Training Objectives
The detection head training loss balances four objectives (complete formulations in Appendix C):
@@ 3.3.2 Detection Head Training Objectives
  • Physics constraints: Penalizes discharge-stage inconsistencies (e.g., deviation from local rating-curve fits) and discourages physically implausible corrections
Training proceeds for 12 epochs with AdamW ($lr=5\times 10^{-5}$), requiring approximately 8 GPU-hours. Because the backbone remains frozen, only the  10K detection head parameters are updated, enabling efficient adaptation while preserving the pretrained backbone representations.
@@ 3.4 Inference Pipeline
At deployment, the trained model processes incoming observations through the backbone to generate reconstructions, then computes detection features from observation-reconstruction residuals (Figure 3, bottom). The inference pipeline produces three outputs per timestep:
  • Anomaly probability: A score in [0, 1] indicating likelihood of data quality issues
  • Uncertainty estimate: Standard deviation across MC Dropout passes (N=20), flagging low-confidence predictions for mandatory review
  • Suggested corrections: Reconstructed discharge ($\hat{Q}$) and stage ($\hat{H}$) values that require hydrologist approval before integration into official records
A three-tier decision system operationalizes these outputs: (1) high-confidence clean predictions (low anomaly probability, low uncertainty) pass with minimal review; (2) high-confidence anomaly detections trigger automatic flagging with suggested corrections; (3) uncertain predictions (high uncertainty regardless of probability) require mandatory expert inspection. Thresholds for the three tiers are selected on the validation set and reported in Section 5. This human-in-the-loop design ensures data quality while enabling continuous monitoring across thousands of stations.
@@ 4.1 Computational Infrastructure
All experiments were run on the Lonestar6 supercomputer at the Texas Advanced Computing Center, accessed through the NSF-led National AI Research Resource (NAIRR) Pilot program [77, 49]. Stage 1 pretraining ran for 20 epochs and required approximately 48 GPU-hours on a single NVIDIA A100 GPU; Stage 2 detection head training ran for 12 epochs on a single NVIDIA A100 GPU. We used mixed-precision training with PyTorch and deterministic seeds fixed per experiment. Unless otherwise stated, we select checkpoints using the lowest validation loss (from clean validation sequences only) on the USGS validation partition (n=798 sites; Section 2.1.4), and we never use the USGS test partition or any Canadian station for model selection.
@@ 4.2 Evaluation Metrics
We evaluate HydroGEM along two complementary dimensions: anomaly detection performance and reconstruction quality.
For anomaly detection we report precision, recall, and F1 score. F1 is our primary reporting metric because it matches the operational setting, where false alarms and missed anomalies both carry cost under a fixed decision rule. Unless otherwise stated, binary predictions are obtained by thresholding anomaly probabilities at 0.5, which we use as a fixed default threshold for all experiments to avoid post hoc tuning.
@@
$\text{Error Reduction}=\frac{\lvert X^{(\text{raw})}-X^{(\text{clean})}\rvert-\lvert\hat{X}-X^{(\text{clean})}\rvert}{\lvert X^{(\text{raw})}-X^{(\text{clean})}\rvert+\epsilon}\times 100\%,$ (1)
where $X\in\{Q,H\}$ and $\epsilon$ is a small constant for numerical stability. We compute metrics separately for discharge and stage and report segment-level averages together with root mean squared error (RMSE) on anomalous segments. We additionally track RMSE on non-anomalous timesteps to confirm that the model does not degrade clean observations. Error reduction and RMSE are computed after denormalization in physical units, restricted to injected anomalous intervals for anomalous-segment metrics.
For the Canadian case study, reconstructions are evaluated primarily as a diagnostic. Weak labels are derived from operational corrections as described in Section 2.3.2, and correction strategies differ between USGS and Environment and Climate Change Canada, particularly for ice-affected periods [54, 58, 20]. As a result, we treat reconstruction error as a secondary diagnostic rather than a primary performance metric and place emphasis on detection metrics.
Because the Canadian weak labels are derived from operational corrections recorded with daily granularity rather than precise anomaly boundaries, we report two complementary detection metrics. First, we report pointwise F1, where predictions and labels are compared at each hourly timestep. Second, we report tolerant F1 with a $\pm$24-hour buffer, which credits a prediction if it falls within $\pm 24$ hourly timesteps of a labeled correction timestep. In addition, we report segment-level recall to evaluate event detection under boundary uncertainty: contiguous labeled correction intervals are treated as anomaly events and an event is counted as detected if any predicted anomaly overlaps the interval or falls within the tolerance window. This multi-metric approach follows recommendations for evaluating time-series anomaly detection under label uncertainty [33, 76].
@@ 4.3 Baseline Methods
To assess zero-shot anomaly detection performance we compare HydroGEM against 11 baseline methods that do not require labeled training data or site-specific calibration. Specifically, we include 3 statistical baselines (Z-Score, IQR, Moving Average Residual), 3 generic unsupervised baselines (Isolation Forest, LOF, STL residual), and 5 hydrology-motivated baselines (rating-curve residual, rate-of-change limits, persistence detection, QQHH consistency checks, and a seasonal envelope). All baselines operate in a zero-shot mode at new stations, matching HydroGEM’s deployment setting. Appendix J contains full baseline parameters and implementation details. For fair comparison, all baselines operate on the same hourly discharge and stage series and use fixed hyperparameters shared across all stations. We consider three categories.
@@
These methods represent simple distributional checks that are widely used in practice. They include a Z-Score detector that flags observations with absolute z-score greater than 3 for either discharge or stage, an Interquartile Range (IQR) rule that flags values outside the interval $Q_{1}-1.5\times\text{IQR},Q_{3}+1.5\times\text{IQR}]$, and a Moving Average Residual detector that uses a centered 7-day window and flags points whose residuals exceed 3 standard deviations of the rolling window. Together these capture pointwise outliers and deviations from local temporal context.
@@
We include generic unsupervised anomaly detectors that require no labels and can be applied at scale. Isolation Forest and Local Outlier Factor (LOF) operate on standardized discharge and stage and use fixed contamination and neighborhood settings specified in Appendix J, representing tree-based and density-based approaches to unsupervised outlier detection. A Seasonal Trend decomposition using LOESS (STL) baseline models a 7-day seasonal component and flags residuals that exceed 3 standard deviations. These methods provide a representative sample of modern off-the-shelf anomaly detectors that do not use hydrological structure.
@@
The third group encodes relationships that hydrologists routinely use during manual quality control. A Rating Curve Residual baseline fits a power-law relationship $Q=a(H-H_{0})^{b}$ by log-linear regression and flags points with large log-space residuals. This serves as a key reference because rating residuals are central to operational practice for detecting rating shifts, drift, and ice effects. Additional checks include rate-of-change limits on discharge and stage, persistence detection for stuck sensors, QQHH consistency checks that flag periods where discharge and stage trends diverge or correlations break down, and a seasonal envelope that flags values outside month-specific percentile bands. Together these methods approximate the mix of physical rules and heuristics that agencies use in production systems.
@@
We intentionally exclude methods that rely on extensive labeled anomalies or site-specific tuning. Supervised deep learning classifiers and autoencoders require labels for each station and would not represent a true zero-shot setting. Per-site reconstruction models that are fit directly on test windows are also excluded, since they would not be deployed in this way operationally. Proprietary automated flagging systems from agencies such as USGS and Environment and Climate Change Canada are not reproducible without internal configuration parameters. Pretraining multiple alternative deep architectures on the full USGS corpus would be methodologically interesting but is beyond the present computational scope and is left for future work.